# Operation of silicon photomultipliers in a dilution refrigerator down to 9.4 mK towards a cryogenic cosmic ray muon veto system

### v1

@@ Abstract
We report the characterisation of a FBK NUV-HD-cryo silicon photomultiplier (SiPM) sensor operated in an 9.4 $\pm$ 0.2 mK environment inside a dilution refrigerator, towards the development of a cryogenic cosmic ray muon veto system to be operated internal to a dilution refrigerator required for low background experiments such as the QUEST-DMC dark matter search experiment. We characterise the single photon response and the gain (the charge produced per detected photon), the dark count noise rate, and correlated noise contributions as a function of operating voltage. This paper also reports first proof of concept measurements of using a SiPM directly coupled to scintillator in a 9.4 mK temperature environment, towards detecting candidate cosmic ray muon signals.
@@ 1 Introduction
In recent years, silicon photomultipliers (SiPMs) have emerged as promising single photon detectors for use in particle physics, particularly in low background, rare-event search experiments such as direct dark matter detection experiments. Photomultiplier Tubes (PMTs) have long been the favoured technology for dark matter experiments, such as DEAP-3600 and LZ [DEAPPMT, LZPMT], however compared to PMTs, SiPMs are cheaper to manufacture, require a lower operational voltage, and are insensitive to magnetic fields. In addition to their high gain and high photon detection efficiency, they can have a lower intrinsic radioactivity compared to PMTs. One drawback of SiPMs is that they generally exhibit higher intrinsic thermal noise than PMTs. However, since thermal noise decreases proportionally with temperature, it can be suppressed by operating the SiPMs at cryogenic temperatures.
@@
Whilst limited, there have been measurements showing that SiPMs can operate close to liquid helium (LHe) temperatures and below [SiPM4K, hanski2025performance, zhang2022scintillation], illustrating that there is a growing interest to demonstrate the viability of operating these sensors in sub-Kelvin conditions. As such, SiPMs are being explored by the QUEST-DMC collaboration as a potential photon sensor for use in a cryogenic cosmic ray muon veto system. The QUEST-DMC collaboration is developing, in one of its work packages, a direct dark matter search experiment projected to reach world-leading sensitivity to spin-dependent sub-GeV/c2 mass dark matter using a superfluid helium-3 (3He) target enclosed in a $\sim$1 cm3 bolometer box [QUESTsensitivity, QUEST-DMC:2025qsa, QUEST-DMC:2025miz]. The QUEST-DMC collaboration is operating bolometers at Royal Holloway, University of London and Lancaster University, in ultra-low temperature dilution refrigerators fitted with an adiabatic demagnetisation stage capable of cooling 3He to operating temperatures below 300 µK.
@@
Since QUEST-DMC is currently operated at sea level, and not underground like the majority of dark matter experiments, cosmic ray induced background interactions will be a limiting factor for QUEST-DMC’s dark matter search [autti2024quest]. The cosmic ray muon veto system is currently in development; the baseline design makes use of a scintillator volume surrounding the bolometer cell at $\sim$300 µK, where the exterior of this volume is coupled to the inner vacuum can which itself is coupled to the 10 mK stage of the dilution refrigerator, and is where the SiPM will be mounted. Therefore, 10 mK is the relevant temperature to characterise the SiPM in order to demonstrate the feasibility of this scheme. We are considering SiPMs over technologies traditionally deployed for photon detection in this temperature range due to the fact that they are large-area, commercially manufacturable, and simple to operate, compared with Superconducting Nanowire Single-Photon detectors (SNSPDs) which are much smaller, and Transition Edge Sensors (TESs) which have a relatively complex fabrication process.
[…]
@@
In this paper, we show the first operation of a NUV-HD-cryo SiPM [NUV-HD-cryo] in a dilution refrigerator, demonstrating the feasibility of deploying this technology in a cryogenic cosmic ray muon system for the QUEST-DMC experiment. The SiPM is attached to the mixing chamber plate of the refrigerator, reaching a base temperature of 9.4 $\pm$ 0.2 mK. Section 2 outlines the experimental setup, the data acquisition, and data processing chain. Section 3 describes the characterisation analysis that was performed. Section 4 reports first proof-of-principle measurements of using a SiPM directly coupled to a scintillator, and operated inside a dilution refrigerator.
@@ 2.1 Hardware
We operate a single NUV-HD-cryo SiPM [NUV-HD-cryo] of 12 $\times$ 8 mm2 inside a cryogen-free dilution refrigerator located at Royal Holloway, University of London. The dilution refrigerator is a modified Oxford Instruments Triton 200 [triton]. A simplified schematic of the experimental setup is shown in Figure 1. The SiPM is mounted onto a custom PCB that distributes bias voltage and housed inside a 7 cm (L) x 3.8 cm (W) x 1.4 cm (D) copper box, which is attached to a copper stand that is thermally coupled to the plate of the mixing chamber.
@@
As Figure 1 shows, all associated electronics are located outside of the dilution refrigerator at room temperature, including a custom-designed signal amplification board, providing a 130 V/V voltage gain. A Keysight E3649A provides the low-voltage supply required to power the board. A Keithley 2450 sourcemeter is used to deliver the bias voltage to the SiPM via the custom amplifier board and an external coaxial cable that connects to the top of the cryostat; an internal coaxial cable then delivers the bias voltage down to the SiPM mounted on the mixing chamber plate on the cold side. In addition, an LED that emits in the red wavelength range is attached to the outside of the copper box, where there is a 1 mm clearance hole allowing a direct line of sight from the LED to the SiPM. The LED is connected to a signal pulse generator via a current-limiting resistor, located outside of the dilution refrigerator, which is used to flash the LED for specific measurements, discussed in Section 2.2. Data is acquired with a Tetronix DPO72304DX 23 GHz bandwidth oscilloscope [tetronixscope].
@@
Figure 2: Left: Picture of the single NUV-HD-cryo SiPM tested in this work, mounted onto a PCB and housed inside a 7 cm (L) x 3.8 cm (W) x 1.4 cm copper box. An LED that emits in the red wavelength range is glued to the outside of the copper box, over a 1 mm clearance hole allowing direct line of sight to the SiPM for acquiring I-V curves in dark conditions. Right: Picture of the copper box (shown in left figure) mounted inside the dilution refrigerator. The copper box is attached to a copper stand thermally coupled to the mixing chamber plate, which reaches a base temperature of 9.4 $\pm$ 0.2 mK.
2.2 I-V curve measurement and breakdown voltage
We first measured the reverse I-V curve of the SiPM using the Keithley 2450 sourcemeter both at room temperature (prior to the cool-down), and after the cool-down to $T_{\mathrm{MC}}$. The reverse I-V curve is used to characterise the relationship between the current flowing through the SiPM and the applied bias voltage, and to determine its breakdown voltage, defined as the minimum bias voltage above which the SiPM can operate in Geiger mode and detect single photons.
At cryogenic temperatures and in a dark environment, an external light source is required to observe the breakdown of the SiPM. A signal pulse generator (1 kHz, 999 µs pulse width) and a current-limiting resistor were used to bias the LED whilst the I-V curve was acquired. The reverse I-V was measured between 0 - 35 V in 300 mV steps; the current limit on the Keithley 2450 was set to 60 µA to avoid damage to the SiPM. Figure  3 (right) shows the measured reverse I-V curves of the SiPM acquired at $T_{\mathrm{MC}}$ (red) and at room temperature prior to the cool-down (green). We observed that whilst acquiring the I-V curve at $T_{\mathrm{MC}}$, the temperature of the mixing chamber plate increased to $\sim$80 mK as a result of flashing the LED for approximately 15 seconds. After the reverse I-V curve was acquired and the LED was switched off, the temperature of the mixing chamber plate returned to $T_{\mathrm{MC}}$ within no more than 2 hours. We note that it is unlikely the temperature of the SiPM actually increased during the reverse I-V measurement, given the short duration (15 seconds) of heat injection and the low thermal conductivity of the SiPM PCB.
Figure 3: Left: Measured reverse I-V curve of the SiPM at $T_{\mathrm{MC}}$ (magenta) and room temperature (green), with breakdown voltages $V_{\mathrm{bd,MC}}$ = 27.16 $\pm$ 0.05 V ($V_{\mathrm{bd,RT}}$ = 32.18 $\pm$ 0.07 V) indicated by dashed red (green) lines. Right: Derivative of the reverse I-V curve, $\mathrm{dI/dV}$, as a function of $V_{\mathrm{bias}}$, acquired at $T_{\mathrm{MC}}$. $V_{\mathrm{bd}}$ is defined as the voltage at which $\mathrm{dI/dV}$ crosses a threshold of $\mu+3\sigma$, indicated by the blue dashed line.
The breakdown voltage, $V_{\mathrm{bd}}$, can be calculated using the reverse I-V curve. The value of $V_{\mathrm{bd}}$ is temperature-dependent, and decreases as a function of lower temperatures [NUV-HD-cryo]. However, it is expected that $V_{\mathrm{bd}}$ will eventually plateau and reach a minimum at a given temperature, such that there is no further reduction in $V_{\mathrm{bd}}$ for temperatures below this value. Our method to calculate $V_{\mathrm{bd}}$ is as follows. First, we calculate the derivative of the I-V curve, $\mathrm{dI/dV}$, and then calculate the mean ($\mu$) and standard deviation ($\sigma$) of $\mathrm{dI/dV}$ between 20 V $\leq V_{\mathrm{bias}}\leq$ 25 V. $V_{\mathrm{bd}}$ is defined as the voltage at which $\mathrm{dI/dV}$ crosses a threshold of $\mu+3\sigma$. Figure 3 (right) shows $\mathrm{dI/dV}$ as a function of $V_{\mathrm{bias}}$, calculated from the I-V curve acquired at $T_{\mathrm{MC}}$, with the threshold indicated by the horizontal red line.
The breakdown voltage of the SiPM at $T_{\mathrm{MC}}$ is calculated to be $V_{\mathrm{bd,MC}}$ = 27.16 $\pm$ 0.05 V, and at room temperature is $V_{\mathrm{bd,RT}}$ = 32.18 $\pm$ 0.07 V. These are indicated by the dashed red (green) lines on Figure 3 (left). As expected, the breakdown voltage decreases with lower temperatures. As the selected photon sensors for the upcoming DarkSide-20k dark matter experiment, the collaboration has tested 359040 NUV-HD-cryo SiPMs at 77 K, measuring several SiPM characteristics including the breakdown voltage. The DarkSide-20k collaboration determined the average breakdown voltage at 77 K to be $\bar{V}_{\mathrm{bd}}$ = 27.19 $\pm$ 0.05 V [QAQC]. This result is consistent with our measurement at $T_{\mathrm{MC}}$, which indicates that there is no further reduction in $V_{\mathrm{bd}}$ for temperatures below 77 K for the NUV-HD-cryo technology.
We report all characterisation measurements in the following subsections as a function of both the operating bias voltage, $V_{\mathrm{bias}}$, and the overvoltage, $\Delta V$, defined as $\Delta V$ = $V_{\mathrm{bias}}-V_{\mathrm{bd}}$.
@@ 2.3 Data acquisition and processing
We acquired three long datasets corresponding to three different operating bias voltages: $V_{\mathrm{bias}}$ = 32, 33, 34 V (equivalently $\Delta V$ = 4.8, 5.8, 6.8 V). Each dataset was acquired in dark conditions, i.e., the LED was switched off. In principle, there should be no other external light reaching the SiPM, as this fridge is inherently dark due to the significant radiation shielding implemented to enable reaching ultra-low temperatures. Therefore, the only signals we expect to observe in these datasets should originate from thermally-induced intrinsic noise associated with the SiPM known as the dark count, plus any additional correlated noise sources. The dark count rate of the NUV-HD-cryo technology is expected to be extremely low, as it is measured to be <<5 mHz/mm2 at 77 K [NUV-HD-cryo].
@@
Approximately 30,000 waveforms were acquired in self-trigger mode per voltage. Data were acquired using the Tetronix FastFrame Memory scheme [FastFrame], which segments the available memory in the oscilloscope such that each segment can be filled with a triggered acquisition at the desired sample rate with minimal dead time between the acquisitions. This is ideal for applications where pulses are separated by large intervals of inactivity in the signal [FastFrame], such as in our case. We acquire waveforms with a 250 MS/s sampling rate and a 100 µs acquisition window, requiring a 5 µs pre-trigger window. We set a 5 mV trigger level on the oscilloscope, which is terminated at 50 $\Omega$, to trigger on the rising edge of the pulse. The trigger level was adjusted by eye in order to maximise triggers from real pulses and minimise triggers from electronics noise, and corresponds to approximately 90% of the mean height of a single photon pulse at $V_{\mathrm{bias}}$ = 32 V. Data is stored in a binary file format, saving the waveform data for each trigger, as well as the trigger time.
@@
Each dataset is analysed using a simple pulse finding algorithm, in order to identify all candidate pulses in each 100 µs-long waveform, and not just the pulse which triggered the event. This ensures that we accurately identify all single photon signals, some of which may have been missed during acquisition due to the 5 mV trigger level that was chosen to eliminate triggers from noise. In each 100 µs-long waveform, a moving average filter is first applied to smooth out noise fluctuations. A baseline-subtracted waveform is computed using the baseline mean calculated in the first 3.2 µs of the waveform, in the pre-trigger region. We then search for regions of the waveform which cross a given threshold for a certain consecutive number of samples, which can be identified as a candidate pulse. The threshold is set to 60% of the filtered mean height of a single photon pulse. We set a looser threshold in the pulse finding algorithm to ensure that we do not miss pulses with amplitudes <<1 PE, which could occur from correlated delayed avalanches that occur within the SiPM recharge time. The time of the pulse, $t_{p}$, is defined as the time corresponding to the maximum amplitude of the waveform in the time-over-threshold region. A “prompt” pulse charge is calculated by integrating the waveform between $t_{p}-100$ ns : $t_{p}+250$ ns]. Sections 3.3 and 3.4 discuss in more detail why we choose to calculate a prompt pulse charge, rather than a full pulse charge.
@@
Figure 4 shows an example waveform acquired at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V), corresponding to a single dark count rate signal, which is equivalent to the single photon signal. The raw waveform (blue), the baseline-subtracted and filtered waveform (orange), the pulse finding threshold (red), and the identified pulse time (green) are shown. In addition to the pulse time, pulse amplitude, and pulse charge, we determine the absolute time difference between the current pulse and the preceeding pulse, $dt$, calculated using the time of the pulses within their respective waveforms, the unix timestamp of the waveforms, and the fraction of a second past the unix timestamps of the waveforms. All of this information is then stored in a data file upon which the downstream characterisation analysis described in Section 3 is performed.
@@
Figure 4: Example waveform acquired at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V), corresponding to a single dark count rate signal (equivalent to the single photon signal). Raw waveform (blue), baseline-subtracted and filtered waveform (orange), pulse finding threshold (red), and identified pulse time (green) are shown.

### v2

# Characterisation of silicon photomultipliers in a dilution refrigerator down to 9.4 mK towards a cryogenic cosmic-ray muon veto system

@@ Abstract
We report the characterisation of a FBK NUV-HD-cryo silicon photomultiplier (SiPM) sensor operated in a 9.4 $\pm$ 0.2 mK environment inside a dilution refrigerator, towards the development of a cryogenic cosmic ray muon veto system to be operated internal to a dilution refrigerator required for low background experiments such as the QUEST-DMC dark matter search experiment. We characterise the single photon response and the gain (the charge produced per detected photon), the dark count noise rate, and correlated noise contributions as a function of operating voltage. This paper also reports first proof-of-concept measurements of using a SiPM coupled to scintillator internal to a dilution refrigerator, towards detecting high-energy events consistent with candidate cosmic-ray muon signals.
@@ 1 Introduction
In recent years, silicon photomultipliers (SiPMs) have emerged as promising single photon detectors for use in particle physics, particularly in low background, rare-event search experiments such as direct dark matter detection experiments. Photomultiplier Tubes (PMTs) have long been the favoured technology for dark matter experiments, such as DEAP-3600 and LZ [8, 7], however compared to PMTs, SiPMs are cheaper to manufacture, require a lower operational voltage, and are insensitive to magnetic fields [16]. In addition to their high gain and high photon detection efficiency, they can have a lower intrinsic radioactivity compared to PMTs. One drawback of SiPMs is that they generally exhibit higher intrinsic thermal noise than PMTs. However, since thermal noise decreases proportionally with temperature, it can be suppressed by operating the SiPMs at cryogenic temperatures.
@@
Whilst limited, there have been measurements showing that SiPMs can operate close to liquid helium (LHe) temperatures and below [11, 18, 27], illustrating that there is a growing interest to demonstrate the viability of operating these sensors in sub-Kelvin conditions. As such, SiPMs are being explored by the QUEST-DMC collaboration as a potential photon sensor for use in a cryogenic cosmic-ray muon veto system. The QUEST-DMC collaboration is developing, in one of its work packages, a direct dark matter search experiment projected to reach world-leading sensitivity to spin-dependent sub-GeV/c2 mass dark matter using a superfluid helium-3 (3He) target enclosed in a $\sim$1 cm3 bolometer box [9, 14, 13]. The QUEST-DMC collaboration is operating bolometers at Royal Holloway, University of London and Lancaster University, in ultra-low temperature dilution refrigerators fitted with an adiabatic demagnetisation stage capable of cooling 3He to operating temperatures below 300 µK.
@@
Since QUEST-DMC is currently operated at sea level, and not underground like the majority of dark matter experiments, cosmic-ray induced background interactions will be a limiting factor for QUEST-DMC’s dark matter search [10]. The cosmic-ray muon veto system is currently in development; the baseline design makes use of a scintillator volume surrounding the bolometer cell, where the exterior of this volume is coupled to the inner vacuum can which itself is coupled to the 10 mK stage of the dilution refrigerator, and is where the SiPM will be mounted. Therefore, 10 mK is the relevant temperature to characterise the SiPM in order to demonstrate the feasibility of this scheme. We are considering SiPMs over technologies traditionally deployed for photon detection in this temperature range due to the fact that they are large-area, commercially manufacturable, and simple to operate, compared with Superconducting Nanowire Single-Photon detectors (SNSPDs) which are much smaller, and Transition Edge Sensors (TESs) which have a relatively complex fabrication process.
SiPM technology has previously been used in cryogenic cosmic-ray muon veto systems, such as the NUCLEUS experiment [15], which makes use of a plastic scintillator deployed in a dilution refrigerator at 850 mK. In this configuration, the plastic scintillator is instrumented with wavelength shifting fibres, and is coupled to a SiPM that is thermally coupled to the 300 K part of the cryostat. However, we are exploring the option of directly coupling SiPMs to the plastic scintillator itself, which would require the SiPM technology to be able to operate in mK environments. Without amplifier electronics and in the absence of light, a single NUV-HD-cryo SiPM with a surface area of 12 $\times$ 8 mm2 dissipates $\mathcal{O}$(pW) of power whilst it is biased, based on the device’s properties measured at 77 K [17]. At this level of power dissipation, the heat load should be low enough such that a SiPM could be thermally coupled to a $\mathcal{O}$(10’s) mK temperature stage of a dilution refrigerator without disrupting the overall cooling power of the cryostat.
@@
In this paper, we show the first operation of a NUV-HD-cryo SiPM [17] in a dilution refrigerator, demonstrating the feasibility of deploying this technology in a cryogenic cosmic-ray muon system for the QUEST-DMC experiment. Section 2 outlines the experimental setup, the data acquisition, and data processing chain. Section 3 describes the characterisation analysis and results. Section 4 reports first proof-of-principle measurements of using a SiPM coupled to a scintillator and operated internal to a dilution refrigerator. To conclude, Section 5 discusses whether, based on the characterisation measurements presented in Section 3, the SiPM tested could meet the requirements of a cryogenic cosmic-ray muon veto system, and how this work informs the next-step R&D for the system.
@@ 2.1 Hardware
We operate a single NUV-HD-cryo SiPM [17] of 12 $\times$ 8 mm2 surface area and 30 µm cell pitch inside a cryogen-free dilution refrigerator located at Royal Holloway, University of London. The dilution refrigerator is a modified Oxford Instruments Triton 200 [19]. A simplified schematic of the experimental setup is shown in Figure 1. The SiPM is mounted onto a custom PCB that distributes bias voltage and is housed inside a 7 cm (L) $\times$ 3.8 cm (W) $\times$ 1.4 cm (D) copper box, shown in Figure 2. The copper box is mounted and thermally coupled to the mixing chamber plate via a copper stand attached perpendicular to the mixing chamber plate, also shown in Figure 2.
@@
We estimate the maximum temperature difference between the SiPM and the mixing chamber temperature measured by the CSNT thermometer assuming the worst possible thermal link, i.e., the SiPM is directly mounted onto the FR-4 PCB which itself is sat directly on the mixing chamber plate (SiPM $\rightarrow$ FR-4 $\rightarrow$ mixing chamber). In this scenario, all of the heat would flow vertically through the FR-4 bulk only. First, we estimate the power dissipation of the SiPM whilst operated in the dark, $P_{\mathrm{SiPM}}$, as,
$P_{\mathrm{SiPM}}=\mathrm{DCR}\cdot G\cdot q\cdot V_{\mathrm{bias}},$ (2.1)
where DCR is the dark count rate, $G$ is the gain, $q=1.6\times 10^{-19}$ C is the electron charge, and $V_{\mathrm{bias}}$ is the bias voltage. We take the upper DCR value reported at 77 K from [17] of 5 mHz/mm2. We assume a gain of $3\times 10^{6}$, taken also from [17] for a SiPM with 30 µm pitch and at 5 V of overvoltage. We take $V_{\mathrm{bias}}$ = 34 V to correspond with the highest bias voltage we acquired characterisation data. With these values, we calculate $P_{\mathrm{SiPM}}$ = 8.2 pW, which we round up to 10 pW. We also add an additional 10 pW of power attributed to wiring within the dilution fridge, yielding a total $P_{\mathrm{SiPM}}$ = 20 pW. We calculate the temperature difference between the SiPM and mixing chamber, $\Delta T=P_{\mathrm{SiPM}}\cdot R_{\mathrm{th}}$, where $R_{\mathrm{th}}$ is the thermal resistance between the SiPM and mixing chamber. $R_{\mathrm{th}}$ is defined as:
$R_{\mathrm{th}}=\frac{L}{k\cdot A},$ (2.2)
[…]
As Figure 1 shows, all associated electronics are located outside of the dilution refrigerator at room temperature, including a custom-designed signal amplification board, providing a 130 V/V voltage gain. A Keysight E3649A provides the low-voltage supply required to power the board. A Keithley 2450 sourcemeter is used to deliver the bias voltage to the SiPM via the custom amplifier board and an external coaxial cable that connects to the top of the cryostat; an internal coaxial cable then delivers the bias voltage down to the SiPM mounted on the mixing chamber plate on the cold side. In addition, an LED that emits in the red wavelength range is attached to the outside of the copper box, where there is a 1 mm clearance hole allowing a direct line of sight from the LED to the SiPM. The LED is connected to a signal pulse generator via a current-limiting resistor, both located outside of the dilution refrigerator, which is used to flash the LED whilst acquiring reverse I-V curves, discussed in Section 3.1. Data is acquired with a Tektronix DPO72304DX 23 GHz bandwidth oscilloscope [25].
@@
Figure 2: Left: Picture of the single NUV-HD-cryo SiPM tested in this work, mounted onto a PCB and housed inside a 7 cm (L) $\times$ 3.8 cm (W) $\times$ 1.4 cm copper box. An LED that emits in the red wavelength range is glued to the outside of the copper box using araldite, over a 1 mm clearance hole allowing direct line of sight to the SiPM for acquiring reverse I-V curves in dark conditions. Right: Picture of the copper box (shown in left figure) mounted inside the dilution refrigerator. The copper box is attached to a copper stand thermally coupled to the mixing chamber plate, which reaches a base temperature of $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK.
@@ 2.2 Data acquisition and processing
We acquired three long datasets corresponding to three different operating bias voltages: $V_{\mathrm{bias}}$ = 32 V, 33 V, and 34 V. Each dataset was acquired in dark conditions, i.e., the LED was switched off. In principle, there should be no other external light reaching the SiPM, as this fridge is inherently dark due to the significant radiation shielding implemented to enable reaching ultra-low temperatures. Therefore, the only signals we expect to observe in these datasets should originate from thermally-induced intrinsic noise associated with the SiPM known as the dark count, plus any additional correlated noise sources. The dark count rate of the NUV-HD-cryo technology is expected to be extremely low, as it is measured to be <<5 mHz/mm2 at 77 K [17].
@@
Approximately 300,000 waveforms were acquired in self-trigger mode per voltage. Data were acquired using the Tektronix FastFrame Memory scheme [26], which segments the available memory in the oscilloscope such that each segment can be filled with a triggered acquisition at the desired sample rate with minimal dead time between the acquisitions. This is ideal for applications where pulses are separated by large intervals of inactivity in the signal [26], such as in our case. We acquire waveforms with a 250 MS/s sampling rate and a 100 µs acquisition window, requiring a 5 µs pre-trigger window. We set a 5 mV trigger level on the oscilloscope, which is terminated at 50 $\Omega$, to trigger on the rising edge of the pulse. The trigger level was adjusted by eye in order to maximise triggers from real pulses and minimise triggers from electronics noise, and corresponds to approximately 70% of the raw mean height of a single photon pulse at $V_{\mathrm{bias}}$ = 32 V. Data is stored in a binary file format, saving the waveform data for each trigger, as well as the trigger time.
@@
Each dataset is analysed using a simple pulse finding algorithm, in order to identify all candidate pulses in each 100 µs-long waveform, and not just the pulse which triggered the event. In each 100 µs-long waveform, a moving-average filter with a 80 ns window is applied to smooth out noise fluctuations. A baseline-subtracted waveform is computed using the baseline mean calculated in the first 3.2 µs of the waveform, in the pre-trigger region. We then search for regions of the waveform above a given threshold, which we set to 40% of the filtered mean amplitude of the single photon pulse for that voltage. We record the segment “start” time at which the signal first crosses the threshold, and the segment “end” time at which the signal drops below the threshold. This corresponds to the number of consecutive samples where the threshold condition was satisfied, known as the time-over-threshold (ToT). We set a lower threshold in the pulse finding algorithm to ensure that we are sensitive to pulses with amplitudes less than the single photon pulse, which could occur from correlated delayed avalanches occurring within the SiPM recharge time (discussed further in Section 3.6).
Figure 3 (left) shows the ToT versus amplitude for candidate pulses at $V_{\mathrm{bias}}$ = 32 V. There are three distinct populations which correspond to the 1, 2, and 3 photon pulses. The 1 photon population is slightly asymmetric compared to the 2 and 3 photon populations, extending to lower amplitude and ToT; this likely originates from recharge-limited correlated delayed avalanches. In addition, there is a separate low amplitude-low ToT population, which is likely driven both by recharge-limited correlated delayed avalanches, and noise fluctuations occurring on the tail of a preceding pulse. We decide to apply an additional, but relaxed, requirement on the time-over-threshold (ToT) of $\geq$ 40 ns for pulse finding, to reduce contamination from noise fluctuations.
The consequence of requiring a lower pulse finding threshold is that we have less sensitivity to pulses occurring close together in time; if a secondary pulse occurs before the signal has dropped below the threshold condition from a primary pulse, the pulse finder will not have the sensitivity to resolve the secondary pulse from the primary. Figure 3 (right) shows the ToT projection of the single photon population illustrated by the red box in the left-hand figure. The mean ToT for the single photon pulse is 192 ns, which means that on average, secondary pulses that occur within 192 ns of a primary pulse will not be resolved as a separate pulse. The mean ToT value for the single photon population is consistent to within 20 ns across all three $V_{\mathrm{bias}}$ values we operated at. For cases where a secondary pulse occurs below the signal drops below threshold, the pulse finder will identify the pulse with the higher amplitude, which predominately originates from the primary pulse. This can introduce a small bias in the characterisation of the correlated delayed avalanche probability, characterised in Section 3.6.
Figure 3: Left: Distribution of the time-over-threshold (ToT) versus amplitude for candidate pulses that exceed the pulse finding threshold (40% of the filtered mean amplitude of the single photon pulse) at $V_{\mathrm{bias}}$ = 32 V. Right: ToT projection of the single photon population illustrated by the red box in the left-hand figure.
@@
Figure 4 shows an example waveform acquired at $V_{\mathrm{bias}}$ = 32 V, corresponding to a single dark count rate signal, which is equivalent to the single photon signal. The raw waveform (WF), the baseline-subtracted (BL-sub) and moving-average filtered waveform (MA-filtered WF), the pulse finding threshold, and the identified pulse time (green) are shown. The pulse time, $t_{\mathrm{p}}$, is defined as the time corresponding to the maximum amplitude. We also calculate a pulse area (proportional to charge) by integrating the waveform between $t_{\mathrm{p}}-100$ ns : $t_{\mathrm{p}}+500$ ns]; this is illustrated by the shaded purple region. In addition to this information, we determine the absolute time difference between the current pulse and the preceding pulse, $dt$, calculated using the time of the pulses within their respective waveforms, the unix timestamp of the waveforms, and the fraction of a second past the unix timestamps of the waveforms. All of this information is then stored in a data file upon which the downstream characterisation analysis described in Section 3 is performed.
@@
Figure 4: Example waveform (WF) acquired at $V_{\mathrm{bias}}$ = 32 V, corresponding to a single dark count rate signal, which is equivalent to the single photon signal. Raw WF (blue), baseline-subtracted (BL-sub) and moving-average (MA) filtered WF (orange), pulse finding threshold (red), and identified pulse time (green) are shown. The shaded purple region corresponds to the region $t_{\mathrm{p}}$-100 ns : $t_{\mathrm{p}}$+500 ns] where the pulse integral is computed to measure the charge.